OpenStudy (anonymous):

Find the slope of the line passing through the points (–3, 4) and (2, –1). 1 –1 3/5 - 5/3

OpenStudy (anonymous):

use the slope formula

OpenStudy (anonymous):

i believe it would be D

OpenStudy (anonymous):

don't think so

OpenStudy (anonymous):

y2-y1 / x2-x1 = m -1-4 / 2 + 3 = -5/5 = -1 is m or the slope

OpenStudy (anonymous):

An easy pneumonic for remembering how to calculate the slope is "rise over run." Select one of your points as a starting point, how far above this point is your next point on the y-axis? Relative to your starting point, how far away on the x-axis is your next point? Put these two numbers in a division problem and that's your slope!

OpenStudy (anonymous):

ok so i figured it out the answer woul be B -1

OpenStudy (alekos):

Nicely explained IParke, but it's called a "mnemonic"